# Operationalizing content moderation “accuracy” in the Digital Services Act

### v3

@@ Abstract.
The Digital Services Act [^1] Regulation (EU) 2022/2065 of the European Parliament and of the Council of 19 October 2022 on a Single Market For Digital Services and amending Directive 2000/31/EC (DSA)., recently adopted by the EU, requires social media platforms to report the “accuracy” of their automated content moderation systems. The colloquial term is vague, or open-textured—the literal accuracy (number of correct predictions divided by the total) is not suitable for problems with large class imbalance, and the ground truth and dataset to measure accuracy against is unspecified. Without further specification, the regulatory requirement allows for deficient reporting. In this interdisciplinary work, we operationalize “accuracy” reporting by refining legal concepts and relating them to technical implementation. We start by elucidating the legislative purpose of the Act to legally justify an interpretation of “accuracy” as precision and recall. These metrics remain informative in class imbalanced settings, and reflect the proportional balancing of Fundamental Rights of the EU Charter. We then focus on the estimation of recall, as its naive estimation can incur extremely high annotation costs and disproportionately interfere with the platform’s right to conduct business. Through a simulation study, we show that recall can be efficiently estimated using stratified sampling with trained classifiers, and provide concrete recommendations for its application. Finally, we present a case study of recall reporting for a subset of Reddit under the Act. Based on the language in the Act, we identify a number of ways recall could be reported due to underspecification. We report on one possibility using our improved estimator, and discuss the implications and need for legal clarification.
Content moderation, EU law, Digital Services Act, inferential statistics, variance reduction.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Lawccs: Mathematics of computing Hypothesis testing and confidence interval computation
1. Introduction
@@
Many concerns have been raised about the role of social media in spreading harmful content such as online hate (Siegel, 2020), and these concerns persist despite the fact that social media companies voluntarily moderate their content (Gillespie, 2018). On the other hand, content moderation raises the question of whether companies are interfering with users’ freedom of speech. Research continues to demonstrate that voluntary moderation is inadequate, and these issues will remain unresolved so long as social media companies have conflicting interests in prioritizing platform growth and profit over providing safe online spaces (Pelley, 2021).
@@
In light of these conflicting interests, legislators have increasingly sought to enact legal requirements for content moderation (Langvardt, 2017). Regulating content moderation is a balancing act between conflicting rights: on one hand there is a need to protect the right to the freedom of speech, and on the other to curtail societal harms (Douek, 2021a). This balance must be made all while considering the technical feasibility of such legal requirements, so as to respect companies’ right to conduct business. Transparency reporting obligations, which require companies to release reports about their content moderation systems, are the first and necessary step in the regulatory landscape. Like any legal instrument, transparency requirements must respect the rights of all entities involved, and be defined with care to ensure their relevance.
@@
The Digital Services Act, recently adopted by the EU, is a milestone legal instrument addressing content moderation. The Act promises to have a far-reaching “Brussels Effect” beyond the EU (see §3) and set the standard for other legislation worldwide. It includes a transparency obligation for social media platforms to report the “accuracy” of their content moderation systems. However, the colloquial term “accuracy” is underspecified, and the regulatory requirement, as it stands, allows for deficient reporting. This is the departure point of our work: we operationalize “accuracy” reporting, and in operationalizing we both refine legal concepts and relate them to technical implementation.
@@
In its entirety, our interdisciplinary work demonstrates that interpreting content moderation “accuracy” in the Digital Services Act as precision and recall is both legally and technically sound. The argumentation here can be useful for future regulatory guidelines or legislative amendment. Our contribution can be viewed from two perspectives. From a legal perspective, we identify the sources of underspecification in the “accuracy” reporting requirement and show that precision and recall are appropriate instantiations that reflect the proportional balancing of Fundamental Rights of the EU Charter. From a technical perspective, we formulate the reporting of recall as unbiased statistical estimation and show how such estimation can be made efficient with stratified sampling. With the improved estimator, we demonstrate the feasibility of recall reporting in a case study on Reddit. Our paper contains three parts:
  1. (1)
@@
    §4 formulates the reporting of recall as an unbiased statistical estimation problem. Since naive estimation can pose an disproportionate burden for companies, we propose the use of stratified sampling. In §5 we derive practical recommendations for applying stratified sampling through a simulation study.
  3. (3)
@@
    §6 presents a case study of a recall reporting requirement for Reddit under the Digital Services Act. Based on the language in the Act, we identify a number of ways recall could be reported due to underspecification. We choose one possibility to report, and discuss the implications and need for legal clarification.
2. Background and related work
Count-based metrics are common.
@@ 2. Background and related work
At the time of writing, many social media companies voluntarily report basic content moderation statistics. Even without direct regulation, companies often provide transparency reports to build public trust in their platform. [^2] https://integrityinstitute.org/news/institute-news/integrity-institute-releases-overview-of-online-social-platform-transparency For instance, Twitter reports a few count-based metrics such as the number of content that Twitter took moderation action against in a given reporting period. [^3] https://transparency.twitter.com/ However, these basic statistics can conflate multiple phenomena. When the number of takedowns of hate speech increases, it is not immediately clear whether it is due to (1) an increase in hate speech, (2) an increase in moderation, or (3) a change in their content moderation policy.
Prevalence is one step away from recall
@@
To the best of our knowledge, only Facebook, Youtube, and Snap publicly provide prevalence metrics. As an example, Facebook reports the prevalence of different categories of content violations over time. Such a prevalence metric would address concerns of (1), reflecting increases in hate speech. While the exact methods to compute such a prevalence are unclear, they note that “we calculate this metric by selecting a sample of content seen on Facebook and then labeling how much of it shouldn’t be there.” [^4] https://about.fb.com/news/2019/05/measuring-prevalence/ Facebook’s own internal studies deem prevalence as a major statistic and industry best practice (Plumb, 2019). With the direct connection between prevalence and moderation thoroughness, we speculate that these large companies may be computing a moderation recall to monitor moderation performance internally, by normalizing prevalence with other statistics, such as number of takedowns (Liu, 2019).
Prevalence is used in academic study as well.
Social media research uses prevalence in online discussions to study the impact of real world events, assuming that prevalence online is a proxy for public opinion (O’Connor et al., 2010). There are many examples of social media research using prevalence estimated from classifier outputs, without human annotations. However, completely automated prevalence classification introduces bias, and for the purposes of legal reporting unbiased estimation should be provided. Park et al. (2022) provides unbiased estimates for the prevalence of macro-norm violating comments (violations of implicit Reddit site-wide rules derived from Chandrasekharan et al. (2018)). They use stratified sampling with two strata, using a classifier for stratification. Our results here complement theirs by providing insights into the application of stratified sampling, and how to improve its efficiency.
Metrics required by regulation.
@@
Almost all major social media platforms provide transparency reports for German users as specified by Germany’s Network Enforcement Act (NetzDG). [^5] Netzwerkdurchsetzungsgesetz vom 1. September 2017 (BGBl. I S. 3352). NetzDG requires companies to periodically provide information and statistics related to the content moderation conducted at these companies (requirements which are succeeded by those in the Digital Services Act). Amongst the required information is also the community guidelines by which companies moderate content, showing that additional regulation can address concern (3). We highlight these reports to demonstrate the potential of regulation to standardize the set of transparency metrics reported on. Our work takes the view that regulation should require the reporting of content moderation recall, a prevalence-based metric. Requiring recall can reflect general moderation effectiveness, and encourages companies to periodically conduct studies of their platform design (Harling et al., 2023).
3. The Digital Services Act
@@
The Digital Services Act (DSA), recently adopted by the EU, will be fully applicable from February 17th, 2024 onwards, and introduces a comprehensive new framework for content moderation. Social media providers with more than 45 million average monthly active recipients in the Union (Art. 33(1) DSA), which currently includes U.S.-established platforms Twitter, Instagram, Youtube, and Facebook, are classified as “very large online platforms” (VLOPs) and will have to comply with the full set of requirements set forth in the Act. [^6] https://ec.europa.eu/commission/presscorner/detail/en/ip_23_2413. This work focuses on the transparency reporting obligations on content moderation introduced by Art. 15(1).
The “Brussels Effect”.
@@
The DSA is a milestone legal instrument targeting content moderation. Hence, our study of the Act may have an impact beyond the EU—a phenomenon termed the “the Brussels effect” (Bradford, 2012). For one, nation-states outside of the EU will likely reference the DSA in considering legislation for their own jurisdictions. As companies may already have the infrastructure to satisfy EU requirements, adopting all or some of the DSA poses a low regulatory burden. Moreover, the DSA (like the GDPR) has a broad territorial scope, and requires companies operating in the EU to apply EU standards to all users, regardless of their nationality (Art. 2(1) DSA). Since the EU is the world’s largest single market, companies often choose to abide by EU regulations instead of forfeiting such business opportunity.
@@
3.1. Scope of content moderation
@@
  • Freedom of expression and of information (Art. 11 CFR) protects online content from censorship and moderating measures. Moderating online posts interferes with this right and requires justification.
  • The right to non-discrimination (Art. 21 CFR) forbids any discrimination based on grounds like sex, race, colour, ethnic or social origin and is especially reflected in the EU framework against the expression of hatred (Zufall et al., 2022). [^14] Usually, national constitutions contain a similar guarantee which is again reflected in national laws, e.g., Criminal Law, that once more constitute ”illegal content” in the sense of the DSA. To prevent discrimination, this right may provide the justification for the interference with freedom of expression.
@@
Balancing out these rights is a task not only performed by the legislator upon drafting regulation like the DSA, but also at the time of its subsequent executive implementation and judicial interpretation (Art. 52(5) CFR). Therefore, these rights continue to be relevant not only for regulatory implementing decisions, but also at the level of technical implementation.
3.2. “Accuracy” as a transparency requirement
Along with balancing out the rights and freedoms in §3.1, any automated content moderation system subject to the DSA needs to equally strike this balance. The EU Charter (Art. 52(1)) demands a legal base (here: the DSA) as any person that is subject to an interference must have the possibility to influence the origin of that interference (here: the obligation to moderate content) through democratic elections. In this way, transparency not only enables democratic control and judicial review, but also offers insight into metrics that measure to which extent democratic legitimacy is carried through to the last step of implementing content moderation.
@@
Art. 15(1)(e) DSA only asks for “indicators of the accuracy and the possible rate of error”. From a technical perspective, the term “accuracy” raises a few questions: if we are estimating accuracy in its literal sense (correct number of predictions divided by the total), to which ground truth are we measuring against? In addition, accuracy is not suitable for problems with large class imbalance, so what metrics would be more appropriate? We identify “accuracy” as an open-textured term, which needs interpretation in novel contexts. Without further interpretation, “accuracy” reporting required by the DSA will exhibit inconsistency, and social media providers may each make their own decisions regarding test data and evaluation metrics. They could choose to report the metric most beneficial for themselves, which would undercut the purpose of the regulation.
@@
Interpreting the notion of “accuracy” beyond its literal wording needs to be in line with the legislative purpose of the Act and requires justification. Future legal adaptations of our analysis depends on this legal justification. We attempt to elucidate and define the notion of “accuracy” in Art. 15(1)(e) of the DSA using a method of systematic interpretation in EU law: inter-instrumental interpretation (Grundmann, 2011). By referencing the definitions and systematic understanding from another instrument of EU law, we can make an attempt to apply it to “accuracy.” We refer to the recent proposal for an Artificial Intelligence Act (AI Act), [^15] COM(2021) 206 final. a parallel development in EU law related to specific high-risk automated systems. Similar to the DSA, the AI Act would apply to U.S.-based providers if they are “placing on the market or putting into service AI systems in the Union” (Art. 2(1)(a) AI Act). While the AI Act is specifically targeted towards high-risk systems, it outlines a general regulatory approach to automated classification systems.
@@
With respect to “accuracy”, Art. 15(1) of the AI Act proposal requires “(high-risk) AI systems to be designed and developed in such a way that they achieve, in the light of their intended purpose, an appropriate level of accuracy […] and perform consistently in those respects throughout their lifecycle.” Similarly to the DSA, the draft only mentions “accuracy” and not other metrics such as precision or recall. Upon further inspection, however, the legislative process reveals that the draft was inspired by the policy guidelines of the High Level Expert Group on AI (HLEG, 2020). In these guidelines, the use of accuracy is intended to reflect the “trustworthiness” of an AI system, but a complementary footnote stated that “accuracy is only one performance metric and it might not be the most appropriate depending on the application,” and that the F1 score, false positives, and false negatives may also be used. Hence, any interpretation of “accuracy” beyond its literal wording in the DSA, should use an “appropriate metric depending on the application”.
3.3. Metrics and their legal implications

### v5

@@ Abstract
The Digital Services Act [^1] Regulation (EU) 2022/2065 of the European Parliament and of the Council of 19 October 2022 on a Single Market For Digital Services and amending Directive 2000/31/EC (DSA)., recently adopted by the EU, requires social media platforms to report the “accuracy” of their automated content moderation systems. The colloquial term is vague, or open-textured—the literal accuracy (number of correct predictions divided by the total) is not suitable for problems with large class imbalance, and the ground truth and dataset to measure accuracy against is unspecified. Without further specification, the regulatory requirement allows for deficient reporting. In this interdisciplinary work, we operationalize “accuracy” reporting by refining legal concepts and relating them to technical implementation. We start by elucidating the legislative purpose of the Act to legally justify an interpretation of “accuracy” as precision and recall. These metrics remain informative in class imbalanced settings, and reflect the proportional balancing of Fundamental Rights of the EU Charter. We then focus on the estimation of recall, as its naive estimation can incur extremely high annotation costs and disproportionately interfere with the platform’s right to conduct business. Through a simulation study, we show that recall can be efficiently estimated using stratified sampling with trained classifiers, and provide concrete recommendations for its application. Finally, we present a case study of recall reporting for a subset of Reddit under the Act. Based on the language in the Act, we identify a number of ways recall could be reported due to underspecification. We report on one possibility using our improved estimator, and discuss the implications and areas for further legal clarification.
1 Introduction
@@
Many concerns have been raised about the role of social media in spreading harmful content such as online hate (Siegel 2020), and these concerns persist despite the fact that social media companies voluntarily moderate their content (Gillespie 2018). On the other hand, content moderation raises the question of whether companies are interfering with users’ freedom of speech. Research continues to demonstrate that voluntary moderation is problematic, and these issues will remain unresolved so long as social media companies have conflicting interests in prioritizing platform growth over providing safe online spaces (Pelley 2021).
@@
In light of these conflicting interests, legislators have increasingly sought to enact legal requirements for content moderation. Regulating content moderation is a balancing act between conflicting rights: on one hand there is a need to protect the right to the freedom of speech, and on the other to curtail societal harms (Douek 2021a). This balance must be made all while considering the technical feasibility of such legal requirements, so as to respect companies’ right to conduct business. Transparency reporting obligations, which require companies to release reports about their content moderation systems, are the first and necessary step in the regulatory landscape. Like any legal instrument, transparency requirements must respect the rights of all entities involved, and be defined with care to ensure their relevance.
@@
The Digital Services Act, recently adopted by the EU, is a milestone legal instrument addressing content moderation. It includes a transparency obligation for social media platforms to report the “accuracy” of their content moderation systems. However, the colloquial “accuracy” is underspecified, and the regulatory requirement, as it stands, allows for deficient reporting. This is the departure point of our work: we operationalize “accuracy” by refining legal concepts and relating them to technical implementation. Our work provides an interpretation of content moderation “accuracy” in the Digital Services Act as precision and recall, which we show is both legally and technically sound. By providing this interpretation, we highlight areas for further legal clarification to ensure meaningful reporting. The argumentation here can be useful for future regulatory guidelines or legislative amendment.
@@
Our interdisciplinary contribution can be viewed from two perspectives. From a legal perspective, we identify the sources of underspecification in the “accuracy” reporting requirement and show that precision and recall are appropriate instantiations which reflect the proportional balancing of Fundamental Rights of the EU Charter. From a technical perspective, we formulate the reporting of recall as unbiased statistical estimation and show how such estimation can be made efficient with stratified sampling. With the improved estimator, we demonstrate the feasibility of recall reporting in a case study on Reddit. Our paper contains three parts:
  1. 1.
@@
    §4 formulates the reporting of recall as an unbiased statistical estimation problem. Since its naive estimation can pose an disproportionate burden for companies, we propose the use of stratified sampling. In §5 we derive practical recommendations for applying stratified sampling through a simulation study.
  3. 3.
@@
    §6 presents a case study of a recall reporting requirement for Reddit under the Digital Services Act. Based on the language in the Act, we identify a number of ways recall could be reported due to underspecification. We choose one possibility to report, and discuss the implications and areas for legal clarification.
2 Background and related work
Count-based metrics
@@ 2 Background and related work
At the time of writing, many social media companies voluntarily report basic content moderation statistics. Even without direct regulation, companies often provide transparency reports to build public trust in their platform. [^2] https://integrityinstitute.org/news/institute-news/integrity-institute-releases-overview-of-online-social-platform-transparency For instance, Twitter reports a few count-based metrics such as the number of posts that Twitter took moderation action against in a given reporting period. [^3] https://transparency.twitter.com/ However, these basic statistics conflate multiple phenomena. When the number of takedowns of hate speech increases, it is not immediately clear whether it is due to an increase in hate speech, an increase in moderation, or a change in their content moderation policy.
Prevalence metrics
@@
To the best of our knowledge, only Facebook, Youtube, and Snap publicly provide prevalence metrics. Facebook reports the prevalence of different categories of content violations over time (e.g. 2 in 1000 posts contained hate speech) [^4] https://transparency.meta.com/policies/improving/prevalence-metric/. As prevalence is a normalized measure (the number of content violations divided by the total amount of content on the platform) it is suitable to track increases in content violations. While their exact methods to compute prevalence are unclear, they note that “we calculate this metric by selecting a sample of content seen on Facebook and then labeling how much of it shouldn’t be there.” [^5] https://about.fb.com/news/2019/05/measuring-prevalence/ Facebook’s own internal studies deem prevalence as a major statistic and industry best practice (Plumb 2019). Research in social media has long used prevalence in online discussions to study the impact of real world events. For example, O’Connor et al. (2010) showed that presidential job approval can be well estimated by the prevalence of tweets mentioning the president which have positive sentiment.
Recall
There is a direct connection between prevalence and recall. By dividing the number of violations acted upon (i.e. the number of takedowns) against the total number of violations (the total number of posts multiplied by violation prevalence, see §4.2), we can calculate the recall of content moderation, or rate at which content violations are identified and acted against. Recall is a measure of moderation “thoroughness”, and we speculate that these large companies may be computing a moderation recall internally to monitor moderation performance, by comparing prevalence with other statistics, such as number of takedowns (Liu 2019).
Metrics required by regulation
@@
Almost all major social media platforms provide transparency reports for German users as specified by Germany’s Network Enforcement Act (NetzDG). [^6] Netzwerkdurchsetzungsgesetz vom 1. September 2017 (BGBl. I S. 3352). NetzDG requires companies to periodically provide information and statistics related to the content moderation conducted at these companies (requirements which are succeeded by those in the Digital Services Act). Amongst the required information is also the community guidelines by which companies moderate content. These reports demonstrate the potential of regulation to standardize the set of transparency metrics reported on. By relating existing transparency requirements to content moderation recall, our work highlights how further legal clarification can ensure meaningful reporting. Recall could reflect general moderation effectiveness, and requiring it would encourage companies to periodically conduct studies of their platform design (Harling, Henesy, and Simmance 2023).
3 The Digital Services Act
@@
The Digital Services Act (DSA), fully applicable since February 17th, 2024, introduces a comprehensive new framework for content moderation. Social media providers with more than 45 million average monthly active recipients in the Union (Art. 33(1) DSA), which currently includes U.S.-established platforms Twitter, Instagram, Youtube, and Facebook, are classified as “very large online platforms” (VLOPs) and will have to comply with the full set of requirements set forth in the Act. [^7] https://ec.europa.eu/commission/presscorner/detail/en/ip˙23˙2413. This work focuses on the transparency reporting obligations on content moderation introduced by Art. 15(1).
Territorial scope
@@
The DSA is a milestone legal instrument targeting content moderation. Hence, our study of the Act may have an impact beyond the EU—a phenomenon termed the “Brussels effect” (Bradford 2012). For one, nation-states outside of the EU will likely reference the DSA in considering legislation for their own jurisdictions. As companies may already have the infrastructure to satisfy EU requirements, adopting all or some of the DSA poses a low regulatory burden. Moreover, the DSA (like the GDPR) has a broad territorial scope, and requires companies operating in the EU to apply EU standards to all users, regardless of their nationality (Art. 2(1) DSA).
@@
3.1 Scope of content moderation
@@
  • Freedom of expression and information (Art. 11 CFR) protects online content from censorship and moderating measures. Moderating online posts interferes with this right and requires justification.
  • The right to non-discrimination (Art. 21 CFR) forbids any discrimination based on grounds like sex, race, colour, ethnic or social origin and is especially reflected in the EU framework against the expression of hatred (Zufall et al. 2022). [^15] Usually, national constitutions contain a similar guarantee which is again reflected in national laws, e.g., Criminal Law, that once more constitute ”illegal content” in the sense of the DSA. To prevent discrimination, this right may provide the justification for the interference with freedom of expression.
@@
Balancing out these rights is a task not only performed by the legislator upon drafting regulation like the DSA, but also at the time of its subsequent executive implementation and judicial interpretation (Art. 52(5) CFR). Any automated content moderation system subject to the DSA that interferes with these rights requires a legal base in the form of law (Art. 52(1) CFR). This requirement reflects the democratic idea that any person who is subject to an interference must have the possibility to influence the origin of that interference through democratic elections. Here, the origin of interference is the content moderation system and the obligation to moderate content. Subsequently enacted laws by the legislation can again then build the legal foundation for limiting and balancing out these fundamental rights. Therefore, these rights continue to be relevant not only for regulatory implementing decisions, but also at the level of technical implementation.
3.2 “Accuracy” as a transparency requirement
Transparency requirements in the DSA not only are important legal instruments for revealing how content moderation systems work, but also for monitoring how they interfere with fundamental rights. Transparency requirements may also reflect the degree to which this balancing of rights has been carried out in favor of one right towards another. Therefore, transparency not only enables democratic control and judicial review, but also offers insight into the extent democratic legitimacy is carried through, to the final step of technical implementation.
@@
Art. 15(1)(e) DSA only asks for “indicators of the accuracy and the possible rate of error”. [^16] The current preparatory document for a draft on a Commission Implementing Regulation laying down templates concerning the transparency reporting obligations (Ares(2023)8428591) does not provide further guidance, and only asks for the reporting of an “accuracy rate” and “error rate” as quantitative information in its templates. From a technical perspective, the term “accuracy” raises a few questions: if we are estimating accuracy in its literal sense (correct number of predictions divided by the total), to which ground truth are we measuring against? In addition, accuracy is not suitable for problems with large class imbalance, so what metrics would be more appropriate? We identify “accuracy” as an open-textured term, which needs interpretation in novel contexts. As it stands, “accuracy” reporting required by the DSA will exhibit inconsistency, and social media providers may each make their own decisions regarding test data and evaluation metrics. They could choose to report the metric most beneficial for themselves, which would undercut the purpose of the regulation. By providing a legally and technically sound interpretation of “accuracy”, our work highlights areas for further legal clarification to avoid an enforcement deficit.
@@
Interpreting the notion of “accuracy” beyond its literal wording needs to be in line with the legislative purpose of the Act and requires justification. Future legal adaptations of our analysis depends on this legal justification. We attempt to elucidate and define the notion of “accuracy” in Art. 15(1)(e) of the DSA using a method of systematic interpretation in EU law: inter-instrumental interpretation (Grundmann 2011). By referencing the definitions and systematic understanding from another instrument of EU law, we can make an attempt to apply it to “accuracy.” We refer to the upcoming Artificial Intelligence Act (AI Act), [^17] Compare the latest consolidated version adopted by the European Parliament: P9_TA(2024)0138. a parallel development in EU law related to specific high-risk AI systems. Similar to the DSA, the AI Act applies to U.S.-based providers if they are “placing on the market or putting into service AI systems in the Union” (Art. 2(1)(a) AI Act). While the AI Act is specifically targeted towards high-risk AI systems, it outlines a general regulatory approach to automated classification systems.
@@
With respect to “accuracy”, Art. 15(1) of the AI Act requires “(high-risk) AI systems to be designed and developed in such a way that they achieve, in the light of their intended purpose, an appropriate level of accuracy […] and perform consistently in those respects throughout their lifecycle.” In this context, the AI Act (Art. 15(2)) calls upon the Commission to encourage “the development of benchmarks and measurement methodologies” for, inter alia, “the appropriate levels of accuracy”. However, similarly to the DSA, the Act only mentions “accuracy” and not other metrics such as precision or recall. Upon further inspection, however, the legislative process reveals that the Act was inspired by the policy guidelines of the High Level Expert Group on AI (HLEG 2020). In these guidelines, the use of accuracy is intended to reflect the “trustworthiness” of an AI system, but a complementary footnote stated that “accuracy is only one performance metric and it might not be the most appropriate depending on the application,” and that the F1 score, false positives, and false negatives may also be used. Hence, any interpretation of “accuracy” beyond its literal wording in the DSA, should use an “appropriate metric depending on the application”.
3.3 Metrics and their legal implications